We test whether “cramming” for an exam is harmful to grades. Condition 1 crams for a pretend exam, but condition 2 does not. Each group has 31 participants.

Below are the data:

M1=43, SS1=1920, M2=48, SS2=2508

a) What is the null and alternative hypothesis?

b) Compute tobt?

c) Compute df?

d) Given the α=0.05, compute the value of tcrit?

e) Is tobt significant?
Answer #1

Ho :   µ1 - µ2 =   0
Ha :   µ1-µ2 ╪   0

b)

difference in sample means =    x̅1-x̅2 =    43.0000   -   48.0   =   -5.000  
                          
pooled std dev , Sp=   √([SS1 + SS2]/(n1+n2-2)) =    8.5907                  
std error , SE =    Sp*√(1/n1+1/n2) =    2.1820                  
                          
t-statistic = ((x̅1-x̅2)-µd)/SE = (   -5.0000   -   0   ) /    2.18   =   -2.2914

c)

      
Degree of freedom, DF=   n1+n2-2 =    60

d) t-critical value , t* = ±2.0003   (excel formula =t.inv(α/2,df)

e) test stat <-2.0003 ,reject Ho

so, t obt is significant